Lehrer

Personalnummer Name  Vorname  Geburtsdatum
0001  Schiller  Friedrich  10.11.1759
0002  Balzac  Honoré de  20.05.1799
0003  Sand  George  01.07.1804
0004  Gauss  Johann Carl Friedrich 30.04.1777
0005  Darvin  Charles Robert  12.02.1809
0006  Curie  Marie  07.11.1867
0007  Friedman  Milton  31.07.1912
0008  Redlich  Oswald  17.09.1858
0009  Dufour  Guillaume-Henri  15.09.1787
0010  Kollwitz  Käthe  08.07.1867
0011  Callas  Maria  02.12.1923
0012  Lovlace  Ada  10.12.1815
0013  Piaget  Jean  09.08.1896
0014  Montessori Maria  31.08.1870